OpenStudy (anonymous):

Find distance between the pair of points: (9,1) and (18,13) Distance in units?

OpenStudy (anonymous):

distance equation sqrt((x2-x1)^2+(y2-y1)^2)

OpenStudy (anonymous):

15

OpenStudy (anonymous):

\[\sqrt{(18-9)^2+(13-1)^2}\]

OpenStudy (anonymous):

\[\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

\[\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

\[\sqrt{9^2+12^2}=\sqrt{81+144}=\sqrt{225}=15\]